# Communication-Efficient Stochastic Distributed Learning

### v1

@@ A-A Step-size bounds
$\kappa_{1}$ $\coloneqq\frac{72\tau}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{2}$
$\kappa_{4}$ $\coloneqq\frac{8}{N}\left((\frac{72\tau}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{2})\right.$
×$\left.\times(\frac{L^{2}}{2}+{2L\rho^{2}d_{u}^{2}\tau+2\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\tau}{|\tilde{\lambda}_{\max}|\rho}(18\rho^{2}d_{u}^{2}+18\rho^{2}d_{u}^{2}\tau L)\right),$
$\tilde{s}_{0}$ $\coloneqq\frac{36m_{u}}{|\tilde{\lambda}_{\max}|\rho}+\frac{16m_{u}}{m_{l}}\left(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)$
$\tilde{s}_{1}$ $\coloneqq\left(\frac{72\tau}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{2}\right)\frac{4m_{u}}{m_{l}}$
$\tilde{s}_{2}$ $\coloneqq\frac{64m_{u}}{m_{l}}\tau^{2}N+\frac{16m_{u}}{m_{l}}N,$
$\tilde{\beta}_{0}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}4(3L^{2}+\rho^{2}d_{u}^{2})$
$+{6L^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}(2\rho^{2}d_{u}^{2}+4L^{2})$
$\tilde{\beta}_{1}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau\rho d_{u}^{2}{72}/{|\tilde{\lambda}_{\max}|}$
$+L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}18\rho d_{u}^{2}\tau{216}/{|\tilde{\lambda}_{\max}|}$
$\tilde{\beta}_{3}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}8L^{2}+{6L^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}4L^{2},$
$\alpha_{0}$ $\coloneqq\frac{L^{2}}{2N}+\frac{2L\rho^{2}d_{u}^{2}\tau+2\rho^{2}d_{u}^{2}+4\tau L^{3}}{N}$
$\alpha_{1}$ $\coloneqq\frac{72\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}$
$\alpha_{5}$ $\coloneqq\left({36\rho d_{u}^{2}\tau^{2}+36\rho d_{u}^{2}\tau^{3}L}\right)/(N|\tilde{\lambda}_{\max}|)$
$\mu$ $\coloneqq 32(\alpha_{5}+\alpha_{0}\alpha_{1}+2(48\tau^{2}L^{2}\alpha_{0}+{4\tau L^{3}}/{N})(\tilde{s}_{0}+\tilde{s}_{1})).$
A-B Preliminary transformation
@@
We start by rewriting the algorithm in a compact form. To this end, we introduce the following auxiliary variables: $\mathbf{Z}=\operatorname{col}\{z_{ij}\}_{i,j\in\mathcal{E}}$, $\mathbf{\Phi}_{k}^{t}=\operatorname{col}\{\phi^{t}_{1,k},\phi_{2,k}^{t},...,\phi_{N,k}^{t}\}$, $G(\Phi_{k}^{t})=\operatorname{col}\{g_{1}(\phi_{1,k}^{t}),g_{2}(\phi_{2,k}^{t}),...,g_{N}(\phi_{N,k}^{t})\}$, $\mathrm{F}(\mathbf{X})=\operatorname{col}\{\frac{1}{N}f_{1}(x_{1}),\frac{1}{N}f_{2}(x_{2}),...,\frac{1}{N}f_{N}(x_{N})\}$. Define $\mathbf{A}=\operatorname{blk\,diag}\{\mathbf{1}_{d_{i}}\}_{i\in\mathcal{V}}\otimes\mathbf{I}_{n}\in\mathbb{R}^{Mn\times Nn},$ where $d_{i}=|\mathcal{N}_{i}|$ is the degree of node $i$, and $M=\sum_{i}|\mathcal{N}_{i}|$. $\mathbf{P}\in\mathbb{R}^{Mn\times Mn}$ is a permutation matrix that swaps $e_{ij}$ with $e_{ji}$. If there is an edge between nodes $i$, $j$, then $A^{T}[i,:]PA[:,j]=1$, otherwise $A^{T}[i,:]PA[:,j]=0$. Therefore $\mathbf{A}^{T}\mathbf{P}\mathbf{A}=\tilde{\mathbf{A}}$ is the adjacency matrix.
@@
$\mathbf{X}_{k+1}=\mathbf{X}_{k}-\gamma\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})+\rho\mathbf{A}^{T}\mathbf{A}\mathbf{\Phi}_{k}^{t}-\mathbf{A}^{T}\mathbf{Z}_{k})$ (12a)
@@
$\mathbf{Y}_{k}=\mathbf{A}^{T}\mathbf{Z}_{k}-\nabla F(\mathbf{\bar{X}}_{k})-\rho\mathbf{D}\mathbf{X}_{k}$ (13)
$\tilde{\mathbf{Y}}_{k}=\mathbf{A}^{T}\mathbf{P}\mathbf{Z}_{k}+\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})-\rho\mathbf{D}\mathbf{X}_{k},$
@@
Multiplying both sides of (12b) by $\mathbf{1}^{T}$, and using the initial condition, we obtain $\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k+1}=\rho\mathbf{1}^{T}\mathbf{D}\mathbf{X}_{k+1}$ for all $k\in\mathbb{N}$. As a consequence $\bar{\mathbf{Y}}_{k}=\mathbf{1}\otimes\frac{1}{N}\mathbf{1}^{T}\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})=\mathbf{1}\otimes\frac{1}{N}\sum_{i}\nabla f_{i}(\bar{x}_{k})$, and (12) can be further rewritten as
$\begin{bmatrix}\mathbf{X}_{k+1}\\ \mathbf{Y}_{k+1}\\ \tilde{\mathbf{Y}}_{k+1}\end{bmatrix}=\begin{bmatrix}\mathbf{I}&\gamma\tau\mathbf{I}&\mathbf{0}\\ \rho\tilde{\mathbf{L}}&\rho\tilde{\mathbf{L}}\gamma\tau+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ \mathbf{0}&-\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}\otimes\mathbf{I}_{n}\begin{bmatrix}\mathbf{X}_{k}\\ \mathbf{Y}_{k}\\ \tilde{\mathbf{Y}}_{k}\end{bmatrix}-\mathbf{h}_{k},$ (14)
@@
$\mathbf{h}_{k}$ $=\large[\gamma\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho D\mathbf{\Phi}_{k}^{t}-\rho\mathbf{D}{\mathbf{X}}_{k});$
$\gamma\rho\tilde{\mathbf{L}}\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\mathbf{\Phi}_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})$
$+\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k});-\nabla F(\bar{\mathbf{X}}_{k+1})+\nabla F(\bar{\mathbf{X}}_{k})\large].$
@@
A-C Deviation from the average
The following lemma illustrates how far the states deviate from the average and will be used later in the proofs of Lemmas 2 and 6.
Lemma 1.
Let Assumption 1 hold, when $\gamma<\frac{2}{\tau|\tilde{\lambda}_{\min}|\rho}$,
$\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\|^{2}\leq\frac{18\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\|\widehat{\mathbf{d}}_{k}\|^{2},\quad\|\bar{\mathbf{Y}}_{k}-\mathbf{Y}_{k}\|^{2}\leq 9\|\widehat{\mathbf{d}}_{k}\|^{2},$ (16)
@@ A-C Deviation from the average
where $\delta=1-{|\tilde{\lambda}_{\max}|\rho\tau\gamma}/{2}<1$.
Proof.
By Assumption 1, graph $\mathcal{G}$ is undirected and connected, hence its Laplacian $-\tilde{\mathbf{L}}$ is symmetric; moreover, it has one zero eigenvalue with eigenvector $\boldsymbol{1}$, with all eigenvalues being positive. Denote by $\widehat{\mathbf{Q}}\in\mathbf{R}^{N\times(N-1)}$ the matrix satisfying $\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}=\mathbf{I}_{N}-\frac{1}{N}\mathbf{11}{}^{T}$, $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}=\mathbf{I}_{N-1}$ and $\mathbf{1}^{T}\widehat{\mathbf{Q}}=0$, $\widehat{\mathbf{Q}}^{T}\mathbf{1}=0$. We have that
@@
where $\mathbf{\Theta}=\begin{bmatrix}\mathbf{I}&\gamma\tau\mathbf{I}&\mathbf{0}\\ \rho\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}&\rho\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\gamma\tau\mathbf{I}+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ \mathbf{0}&\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}$.
The next step is to show that $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$ is negative definite by contradiction. Let $x\in R^{N-1}$ be an arbitrary vector, since $-\tilde{\mathbf{L}}$ is the positive semi-definite Laplacian matrix, the quadratic form $x^{T}\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x=(\widehat{\mathbf{Q}}x)^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x\leq 0$. Moreover, if $(\widehat{\mathbf{Q}}x)^{T}(\tilde{\mathbf{A}}-\mathbf{D})\widehat{\mathbf{Q}}x=0$, we have $\widehat{\mathbf{Q}}x=\boldsymbol{1}$. Now, the properties of $\widehat{\mathbf{Q}}$ imply that $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}x=x=\widehat{\mathbf{Q}}^{T}\boldsymbol{1}=0$. Therefore, for all non-zero vectors $x$, the quadratic form $x^{T}\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x<0$, thus $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$ is a symmetric negative-definite matrix.
@@
$=\begin{bmatrix}\mathbf{I}&\gamma\tau&0\\ \rho\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}&\rho\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}\gamma\tau+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ 0&-\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}.$
We denote $\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}=\operatorname{diag}\{\tilde{\lambda}_{i}\}_{i=2,...,N}$, where $\tilde{\lambda}_{i}<0$ is the eigenvalue of $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$, $\tilde{\lambda}_{\min}=\lambda_{\min}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$, and $\tilde{\lambda}_{\max}=\lambda_{\max}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$. Since each block of $\tilde{\mathbf{\Theta}}$ is a diagonal matrix, there exists a permutation matrix $\mathbf{P}_{0}$ such that $\mathbf{P}_{0}\tilde{\mathbf{\Theta}}\mathbf{P}_{0}^{T}=\mathbf{P}_{0}\boldsymbol{\phi}\mathbf{\Theta}\boldsymbol{\phi}^{T}\mathbf{P}_{0}^{T}=\operatorname{blkdiag}\left\{\mathbf{D}_{i}\right\}_{i=2}^{N},$ where
$\mathbf{D}_{i}=\begin{bmatrix}1&\gamma\tau&0\\ \rho\tilde{\lambda}_{i}&\rho\tilde{\lambda}_{i}\gamma\tau+0.5&-0.5\\ 0&-0.5&0.5\end{bmatrix}.$ (20)
@@
$\mathbf{V}_{i}=\begin{bmatrix}-\gamma\tau&d_{12}&d_{13}\\ 1&d_{22}&d_{23}\\ 1&1&1\end{bmatrix}$ (21)
with $d_{12}=-\gamma\tau+((\gamma\tilde{\lambda_{i}}\rho\tau(\gamma\tilde{\lambda_{i}}\rho\tau+2))^{0.5})/(\tilde{\lambda_{i}}\rho)$, $d_{13}=-\gamma\tau-((\gamma\tilde{\lambda_{i}}\rho\tau(\gamma\tilde{\lambda_{i}}\rho\tau+2))^{0.5})/(\tilde{\lambda_{i}}\rho)$, $d_{22}=\tilde{\lambda_{i}}\rho d_{12}-1$, $d_{23}=\tilde{\lambda_{i}}\rho d_{13}-1$. The nonzero eigenvalues $\lambda$ of $\mathbf{D}_{i}$, $i=2,\ldots,N$, satisfy $2\lambda^{2}+(-2\tilde{\lambda}_{i}\rho\tau\gamma-4)\lambda+\tilde{\lambda}_{i}\rho\tau\gamma+2=0,$ which can be written in the form:
$02\lambda^{2}-2t\lambda+t=$ 0 (22)
where $t=\tilde{\lambda}_{i}\rho\tau\gamma+2$. The modulus of the roots of (22) is $1-\frac{|\tilde{\lambda}|\rho\tau\gamma}{2}$ when $-2<\tilde{\lambda}\rho\tau\gamma<0$. We conclude that we can write $\mathbf{\Theta}=(\mathbf{P}_{0}\boldsymbol{\phi})^{T}\mathbf{V}\boldsymbol{\Delta}\mathbf{V}^{-1}(\mathbf{P}_{0}\boldsymbol{\phi})$ where $\mathbf{V}=\operatorname{blkdiag}\left\{V_{i}\right\}_{i=2}^{N}$ and
@@
Moreover, $\|\mathbf{\Delta}\|=1-{|\tilde{\lambda}_{\max}|\rho\tau\gamma}/{2}$ when
$|\tilde{\lambda}_{\min}|\rho\tau\gamma<2.$ (24)

### v2

@@ 7.1 Step-size bounds
$\bar{\gamma}_{13}$ $\coloneqq\frac{\lambda_{l}}{\lambda_{u}\sqrt{8(\kappa_{0}\tilde{\beta}_{0}+2(\tilde{s}_{0}+\tilde{s}_{1})(\kappa_{1}+32\tau^{2}L^{2}\kappa_{0}))}},$
$\bar{\gamma}_{14}$ $\coloneqq\sqrt[3]{\frac{\lambda_{l}^{2}\beta^{2}}{768L^{2}N\lambda_{u}^{2}\kappa_{4}}},\quad\bar{\gamma}_{15}\coloneqq\sqrt[3]{\frac{\lambda_{l}^{2}\tau}{128\lambda_{u}^{2}\kappa_{4}\kappa_{2}}}$
These bounds depend on the following quantities. $d_{u}=\max\{|\mathcal{N}_{i}|\}_{{i\in\mathcal{V}}}$ denotes the maximum agents’ degree. $\widehat{\mathbf{V}}$ is defined in (25). $\lambda_{u}$ is the largest eigenvalue of the graph $\mathcal{G}$’s Laplacian matrix, $\lambda_{l}$ is the smallest nonzero eigenvalue of graph $\mathcal{G}$’s Laplacian matrix. We denote $m_{u}=\max_{i=1,...,N}m_{i}$ and $m_{l}=\min_{i=1,...,N}m_{i}$, where $m_{i}$ is the number of local data points of agent $i$. Additionally, we have the following definitions, used both in the upper bound above and throughout the convergence analysis:
$\beta_{0}$ $\coloneqq\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2},$
$\tilde{\beta}_{0}$ $\coloneqq\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2},$
$c_{4}$ $\coloneqq\frac{4L^{2}}{N}\left(\frac{72\beta\tau}{\lambda_{l}\rho}+144\tau^{2}\beta^{2}\right)$
$\kappa_{0}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})6\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}+\frac{6L^{2}}{\beta^{2}}2\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2},$
$\kappa_{1}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})4\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}+\frac{6L^{2}}{\beta^{2}}2\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2},$
$\kappa_{2}$ $\coloneqq 16\tau^{3}N\kappa_{0}+2\tilde{s}_{2}(\kappa_{1}+32\tau^{2}L^{2}\kappa_{0}),$
$\tilde{s}_{0}$ $\coloneqq\frac{36\beta\tau^{2}m_{u}}{\lambda_{l}\rho}+\frac{144\tau^{2}m_{u}}{m_{l}}\beta^{2},$
$\tilde{s}_{1}$ $\coloneqq\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\frac{8m_{u}\tau}{m_{l}},$
$\tilde{s}_{2}$ $\coloneqq\frac{16Nm_{u}\tau^{2}}{m_{l}}+\frac{8m_{u}\tau}{m_{l}}16\tau^{3}N,$
$\kappa_{3}$ $\coloneqq 16\tau^{3}N(\frac{L^{2}}{2N}+2\tau L^{3})$
$+2\tilde{s}_{2}(2\tau L^{3}+32\tau^{2}L^{2}(\frac{L^{2}}{2N}+2\tau L^{3})),$
$\kappa_{4}$ $\coloneqq(\frac{L^{2}}{2N}+2\tau L^{3})\tilde{\beta}_{0}$
$+2(\tilde{s}_{0}+\tilde{s}_{1})(2\tau L^{3}+32\tau^{2}L^{2}(\frac{L^{2}}{2N}+2\tau L^{3})),$
where $\frac{1}{\tau\lambda_{u}\rho}\leq\beta<\frac{2}{\tau\lambda_{u}\rho}$.
7.2 Preliminary transformation
@@
We start by rewriting the algorithm in a compact form. To this end, we introduce the following auxiliary variables: $\mathbf{Z}=\operatorname{col}\{z_{ij}\}_{i,j\in\mathcal{E}}$, $\mathbf{\Phi}_{k}^{t}=\operatorname{col}\{\phi^{t}_{1,k},\phi_{2,k}^{t},...,\phi_{N,k}^{t}\}$, $G(\Phi_{k}^{t})=\operatorname{col}\{g_{1}(\phi_{1,k}^{t}),g_{2}(\phi_{2,k}^{t}),...,g_{N}(\phi_{N,k}^{t})\}$, $\mathrm{F}(\mathbf{X})=\operatorname{col}\{f_{1}(x_{1}),f_{2}(x_{2}),...,f_{N}(x_{N})\}$, $F({x})=\frac{1}{N}\sum_{i=1}^{N}f_{i}({x})$. Define $\mathbf{A}=\operatorname{blk\,diag}\{\mathbf{1}_{d_{i}}\}_{i\in\mathcal{V}}\otimes\mathbf{I}_{n}\in\mathbb{R}^{Mn\times Nn},$ where $d_{i}=|\mathcal{N}_{i}|$ is the degree of node $i$, and $M=\sum_{i}|\mathcal{N}_{i}|$. $\mathbf{P}\in\mathbb{R}^{Mn\times Mn}$ is a permutation matrix that swaps $e_{ij}$ with $e_{ji}$. If there is an edge between nodes $i$, $j$, then $A^{T}[i,:]PA[:,j]=1$, otherwise $A^{T}[i,:]PA[:,j]=0$. Therefore $\mathbf{A}^{T}\mathbf{P}\mathbf{A}=\tilde{\mathbf{A}}$ is the adjacency matrix.
@@
$\mathbf{X}_{k+1}=\mathbf{X}_{k}-\sum_{t=0}^{\tau-1}(\gamma G(\mathbf{\Phi}_{k}^{t})+{\beta}(\rho\mathbf{A}^{T}\mathbf{A}\mathbf{X}_{k}-\mathbf{A}^{T}\mathbf{Z}_{k}))$ (12a)
@@
$\mathbf{Y}_{k}=\mathbf{A}^{T}\mathbf{Z}_{k}-\frac{\gamma}{\beta}\nabla F(\mathbf{\bar{X}}_{k})-\rho\mathbf{D}\mathbf{X}_{k}$ (13)
$\tilde{\mathbf{Y}}_{k}=\mathbf{A}^{T}\mathbf{P}\mathbf{Z}_{k}+\frac{\gamma}{\beta}\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})-\rho\mathbf{D}\mathbf{X}_{k},$
@@
Multiplying both sides of (12b) by $\mathbf{1}^{T}$, and using the initial condition, we obtain $\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k+1}=\rho\mathbf{1}^{T}\mathbf{D}\mathbf{X}_{k+1}$ for all $k\in\mathbb{N}$. As a consequence $\bar{\mathbf{Y}}_{k}=\frac{\gamma}{\beta}\mathbf{1}\otimes\frac{1}{N}\mathbf{1}^{T}\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})=\frac{\gamma}{\beta}\mathbf{1}\otimes\frac{1}{N}\sum_{i}\nabla f_{i}(\bar{x}_{k})$, and (12) can be further rewritten as
$\begin{bmatrix}\mathbf{X}_{k+1}\\ \mathbf{Y}_{k+1}\\ \tilde{\mathbf{Y}}_{k+1}\end{bmatrix}=\begin{bmatrix}\mathbf{I}&\beta\tau\mathbf{I}&\mathbf{0}\\ \rho\tilde{\mathbf{L}}&\rho\tilde{\mathbf{L}}\beta\tau+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ \mathbf{0}&-\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}\otimes\mathbf{I}_{n}\begin{bmatrix}\mathbf{X}_{k}\\ \mathbf{Y}_{k}\\ \tilde{\mathbf{Y}}_{k}\end{bmatrix}-\mathbf{h}_{k},$ (14)
@@
$\mathbf{h}_{k}$ $=\large[\gamma\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}));$
$\gamma\rho\tilde{\mathbf{L}}\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}))+\frac{\gamma}{\beta}(\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k}));$
$\frac{\gamma}{\beta}(-\nabla F(\bar{\mathbf{X}}_{k+1})+\nabla F(\bar{\mathbf{X}}_{k}))\large].$
@@
7.3 Deviation from the average
The following lemma illustrates how far the states deviate from the average and will be used later in the proofs of Lemmas 8.4 and 9.12.
Lemma 1
Let Assumption 1 hold, when $\beta<\frac{2}{\tau\lambda_{u}\rho}$,
$\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\|^{2}\leq\frac{18\beta\tau}{\lambda_{l}\rho}\|\widehat{\mathbf{d}}_{k}\|^{2},\quad\|\bar{\mathbf{Y}}_{k}-\mathbf{Y}_{k}\|^{2}\leq 9\|\widehat{\mathbf{d}}_{k}\|^{2},$ (16)
@@ 7.3 Deviation from the average
where $\delta=1-{\lambda_{l}\rho\tau\beta}/{2}<1$, $\widehat{\mathbf{d}}_{k}=\widehat{\mathbf{V}}^{-1}\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k};\widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k};\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}\end{bmatrix}$, $\widehat{\mathbf{Q}}$ and $\widehat{\mathbf{V}}^{-1}$ are matrices used to define the deviation term $\widehat{\mathbf{d}}_{k}$.
Proof 7.3.
By Assumption 1, graph $\mathcal{G}$ is undirected and connected, hence its Laplacian $\tilde{\mathbf{L}}$ is symmetric; moreover, it has one zero eigenvalue with eigenvector $\boldsymbol{1}$, with all eigenvalues being positive. Denote by $\widehat{\mathbf{Q}}\in\mathbf{R}^{N\times(N-1)}$ the matrix satisfying $\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}=\mathbf{I}_{N}-\frac{1}{N}\mathbf{11}{}^{T}$, $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}=\mathbf{I}_{N-1}$ and $\mathbf{1}^{T}\widehat{\mathbf{Q}}=0$, $\widehat{\mathbf{Q}}^{T}\mathbf{1}=0$. We have that
@@
where $\mathbf{\Theta}=\begin{bmatrix}\mathbf{I}&\beta\tau\mathbf{I}&\mathbf{0}\\ \rho\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}&\rho\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\beta\tau\mathbf{I}+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ \mathbf{0}&\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}$.
The next step is to show that $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$ is negative definite by contradiction. Let $x\in R^{N-1}$ be an arbitrary vector, since $\tilde{\mathbf{L}}$ is the positive semi-definite Laplacian matrix, the quadratic form $x^{T}\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x=(\widehat{\mathbf{Q}}x)^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x\leq 0$. Moreover, if $(\widehat{\mathbf{Q}}x)^{T}(\tilde{\mathbf{A}}-\mathbf{D})\widehat{\mathbf{Q}}x=0$, we have $\widehat{\mathbf{Q}}x=\boldsymbol{1}$. Now, the properties of $\widehat{\mathbf{Q}}$ imply that $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}x=x=\widehat{\mathbf{Q}}^{T}\boldsymbol{1}=0$. Therefore, for all non-zero vectors $x$, the quadratic form $x^{T}\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x<0$, thus $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$ is a symmetric negative-definite matrix.
@@
$=\begin{bmatrix}\mathbf{I}&\beta\tau&0\\ \rho\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}&\rho\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}\beta\tau+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ 0&-\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}$.
[…]
$\mathbf{D}_{i}=\begin{bmatrix}1&\beta\tau&0\\ \rho\tilde{\lambda}_{i}&\rho\tilde{\lambda}_{i}\beta\tau+0.5&-0.5\\ 0&-0.5&0.5\end{bmatrix}$. (20)
@@
$\mathbf{V}_{i}=\begin{bmatrix}-\beta\tau&d_{12}&d_{13}\\ 1&d_{22}&d_{23}\\ 1&1&1\end{bmatrix}$ (21)
with $d_{12}=-\beta\tau+((\beta\tilde{\lambda_{i}}\rho\tau(\beta\tilde{\lambda_{i}}\rho\tau+2))^{0.5})/(\tilde{\lambda_{i}}\rho)$, $d_{13}=-\beta\tau-((\beta\tilde{\lambda_{i}}\rho\tau(\beta\tilde{\lambda_{i}}\rho\tau+2))^{0.5})/(\tilde{\lambda_{i}}\rho)$, $d_{22}=\tilde{\lambda_{i}}\rho d_{12}-1$, $d_{23}=\tilde{\lambda_{i}}\rho d_{13}-1$. The nonzero eigenvalues $\lambda$ of $\mathbf{D}_{i}$, $i=2,\ldots,N$, satisfy $2\lambda^{2}+(-2\tilde{\lambda}_{i}\rho\tau\beta-4)\lambda+\tilde{\lambda}_{i}\rho\tau\beta+2=0,$ which can be written in the form:
$2\lambda^{2}-2t\lambda+t=0$ (22)
where $t=\tilde{\lambda}_{i}\rho\tau\beta+2$. The modulus of the roots of (22) is $1-\frac{|\tilde{\lambda}|\rho\tau\beta}{2}$ when $2<\tilde{\lambda}\rho\tau\beta<0$. We conclude that we can write $\mathbf{\Theta}=(\mathbf{P}_{0}\boldsymbol{\phi})^{T}\mathbf{V}\boldsymbol{\Delta}\mathbf{V}^{-1}(\mathbf{P}_{0}\boldsymbol{\phi})$ where $\mathbf{V}=\operatorname{blkdiag}\left\{V_{i}\right\}_{i=2}^{N}$ and
@@
Moreover, $\|\mathbf{\Delta}\|=1-{\lambda_{l}\rho\tau\beta}/{2}$ when
$\lambda_{u}\rho\tau\beta<2.$ (24)